<e> Congratulations! You are hired in a business organization. In this company the market research department that you are the member recommended to manufacture and market a promising new product (x). After extensive surveys, the research department supported the recommendation with the demand function:


Dx = f (p) = 60 – 3 P


Where Dx is quantity demanded at price P during a month. Financial department has brought the cost function and is given by


C(x) = 72 + 6x


Now the team leader of the market research department ordered you to manipulate the following problems.


Express revenue as a function of X


Expresses profit as a function of X



1
Expert's answer
2022-03-08T13:06:02-0500

Solution


Given the demand function is


"D(x)=f(P)=60-3P"


Here, "D" is the quantity demanded at price "P"


And the cost function is


"C(x)=72+6x"



Now we know that Revenue is equal to the number of units sold times the price per unit


Therefore, the Revenue is


"R(x) = x\times P"


"R(x) = Px" Revenue function in terms of ""x""



Now, we know that Profit is calculated by subtracting the cost from the revenue collected. Therefore,


"Profit=R(x)-C(x)\\"


"Profit=Px - (72+6x)"


"Profit=(P-6)x -72"